OpenStudy (anonymous):

Find the difference. (4x2 + 3x + 2) - (2x2 + 2x + 1) A. 2x2 + 5x + 3 B. 2x2 + x + 1 C. 2x2 - x - 1

OpenStudy (anonymous):

B

OpenStudy (anonymous):

subtract each set from each other 4x^2 - 2x^2 = 2 x^2 3x-2x = x 2-1 = 1 answer is b